Use elimination to solve the linear system.
3x-4y=21
4x+2y=6

Answer by checkley79(3341)

3x-4y=21 divide by 2 & add.
4x+2y=6
1.5x-2y=10.5
------------------------
5.5x=16.5
x=16.5/5.5
x=3 ans
3*3-4y=21
9-4y=21
-4y=21-9
-4y=12
y=12/-4
y=-3 ans.
Proof:
4*3+2*-3=6
12-6=6
6=6